# In-context learning of evolving data streams with tabular foundational models

### v1

@@ 5 Future research directions
In the early stages of machine learning, the dominant belief was that the best way to solve a task was to meticulously design a specific dataset and tailor a model exclusively for that task. Over time, it became evident that for certain types of unstructured data, such as image recognition, a general-purpose model could be fine-tuned on specific datasets, yielding effective results. Today, we stand at a point where highly versatile models, such as general-purpose transformers, can be prompted to tackle a wide range of tasks without task-specific training. With this work, we embrace this evolving paradigm. So far, we have identified LTMs as a promising candidate for streaming solutions, warranting further research. The key challenge is not merely the size of the model or processing speed but rather the dynamic interplay between data arrival, training, recovery, and inference. As technology advances, models will naturally become smaller, faster, and more affordable. In the meantime, efforts should focus on designing algorithms that align more closely with the core principles of data stream mining: processing data in its natural order, handling labels that may be correct at the time of request but outdated or even incorrect by the time they become available, minimizing interference among concepts, and so on. Building on this foundation, several promising research directions emerge within this evolving paradigm. The authors encourage the academic community to explore these opportunities, as they hold significant potential for advancing the field. The authors, too, may pursue some of these directions and welcome feedback and collaboration from those interested in contributing to this line of research.
5.1 Developing new LTMs
Since the pre-training stage did not represent the added value of this work, an already existing LTM for batch data was used. However, many benefits could be introduced by scaling LTMs with large-scale drifting tabular datasets. Indeed, it has been shown repeatedly that the construction of modality-specific large-scale datasets allows to significantly increase performance, while using the same model architecture and training procedure from prior work [12]. For instance, recent work filtered large collections of raw tabular data [112] with a mix heuristics and rule-based methods applied at a table, column and row level, to extract a high-quality subset for tabular prediction [113]. With this heavy modality-specific pre-training, this LLM now outperforms traditional batch tree-based methods across unseen domains, relying on a small number of representative examples for few-shot learning without any fine-tuning. This shows the performance bottleneck of previous similar efforts [15, 14, 71] lied in a lack of large-scale tabular pre-training. In this regard, it is important to note that developing state-of-the-art LTMs is still within the computational reach of many AI researchers, opposed to the economically-exclusionary cost of training modern LLMs.
5.2 Hypernetwork-based LTMs
Instead of predicting the target directly, hypernetwork-based LTMs predict a set of concept-specific weights for a target model. This target model then performs classification on tabular data, allowing the LTM to remain independent from the inference stage, which is particularly beneficial in memory- and time-constrained scenarios. Examples of such LTMs include HyperFast [28] and MotherNet [29], both of which generate feed-forward neural networks through a single forward pass. In HyperFast, the weights of the main network are generated layer-by-layer by hypernetwork modules with shared and layer-specific parameters. The first layer applies a random features kernel trick combined with PCA to form a fixed-size, permutation-invariant representation. Layers 2 to $L-1$ employ MLP-based hypernetworks that process intermediate representations, incorporating residual connections and non-linear activations. For the final layer $L$, HyperFast avoids directly predicting the weights, instead averaging per-class embeddings to create a lightweight classification weight matrix. In contrast, MotherNet adapts the TabPFN model to generate weights for a compact feed-forward network with a fixed architecture consisting of two hidden layers of 512 units each. MotherNet introduces an attention layer that compresses all activations into a single dataset embedding, based on a learned query vector [29]. This query token is optimized during training and subsequently fixed, similarly to prefix tuning [114], but is applied solely within a top-layer attention module rather than across the entire transformer. For future research, one can draw upon the traditional hypernetwork architectural paradigms, which are typically categorized as either task-conditioned or data-conditioned. Task-conditioned schemes involve providing task-specific inputs, such as task identity/embedding, hyperparameters, or other cues [115], while data-conditioned schemes adapt based on the characteristics of the input data [116, 117]. While hypernetwork-based LTMs have primarily focused on data-conditioned schemes, exploring task-conditioned schemes may offer better information sharing across multiple tasks [115]. However, the conditions under which to model a problem as data-conditioned or task-conditioned remain unclear, and require further exploration depending on the problem, the availability of data, and the number of tasks [118]. Beyond the dynamic nature of input-dependent contexts, future work can investigate alternative output-based strategies. Hypernetwork-based LTMs have primarily focused on generating weights for the entire target model simultaneously [29, 28], but generating a large number of weights for the target model can result in a larger last layer in the hypernetwork. One potential solution is to introduce multiple heads for weight generation, thereby reducing the number of weights required in the final layer by a factor corresponding to the number of heads [115]. Furthermore, adopting a chunk-wise or component-wise weight generation approach could be advantageous. In chunk-wise approaches, some weights may remain unused, with additional embeddings used to distinguish and produce different chunks, which reduces complexity and improves the scalability of hypernetworks [115]. Component-wise approaches go further, generating separate weights for each layer or channel in the target model, allowing the network to focus on specific features or patterns [117]. It is important to note that, unlike traditional hypernetwork architectures [119], which use task-specific, multi-task hypernetworks to generate large target networks, hypernetwork-based LTMs utilize large, transformer-style networks to generate compact tabular classification models. Moreover, hypernetwork-based LTMs should generate all weights in a single forward pass, rather than generating weights for specific layers (e.g., the final layer) while training other parts of the feature extractor [120, 121].
5.3 Localization strategy
Similar to LLMs, compressing the context for LTMs is essential for ensuring alignment with the context length. This process must account for the fact that different data points exhibit varying levels of potential, with some being more representative or informative than others. Rather than naively selecting a random subsample of the training data as context for the prompt [33, 122], it is more effective to retrieve information that is specifically relevant to the task at hand. Techniques such as compressing the dataset through k-means centroids [79] or employing dataset distillation [80] have been applied. In a streaming setting, however, synopsis techniques, such as histograms, wavelets, sketches are often required to construct geometric and statistical data descriptors as the online phase of clustering-based approaches, capturing the internal structure of the classes [123]. Subsequently, these methods are followed by a offline mining process performed on the stored micro-cluster synopsis whenever a user sends a request. For example, this two-phase learning approach has been used as a temporal extension for clustering feature vectors [124], which was later extended with projection techniques for subspace clustering in high-dimensional data streams [125] and temporal-dependent splitting for long-term operation [126]. Moreover, this synopsis strategy can be utilized to build a density profile of the data, aiding in the discovery of arbitrarily shaped clusters by extending DBSCAN [127]. This includes calculating estimates for the epsilon neighborhood around each data point [128], or for micro-clusters [129]. These methods not only reflect the proximity of data points but also capture their interconnectivity [130]. Additionally, these techniques can be incorporated into drift detection schemes to actively adapt in non-stationary environments. These concept drifts can monitor the performance of the LTM itself, e.g. focusing on prediction errors, such as error trend [131] error distance [132], and different rates of the confusion matrix [133], or via posterior probability estimates indicating how well the classes are separated by the classifier, incrementally testing for differences, e.g using KL divergence [134], or a Page-Hinckley test [135]. Furthermore, instead of relying solely on the LTM or a window of recent data to define drift severity, one can explicitly combine drift adaptation with the context optimization strategy. For example, dynamically maintaining a small set of short-term and long-term prototypes based on error-driven representativeness learning and synchronization-inspired constrained clustering [136]. In clustering-driven context optimization, concept drift can be detected by tracking assignment errors [137] or comparing recent and reference data within each cluster using the univariate k-sample Anderson-Darling test for each principal component of each centroid [138]. Several drift detectors rely on clustering to construct data subspaces for regional density estimation. This includes using neighbor entropy to identify similarities [139], or accumulating density discrepancies in each subspace [140]. These density-based approaches can also be implemented as graphs, such as detecting the origin of drift by updating the k-dist graph based on DBSCAN [141].
5.4 Retrieval-based strategy
Context optimization should not be constrained to a uniform context across all query points. While certain coresets may provide valuable insights for defining the current concept, others could limit the model’s adaptability. Instead, a more flexible approach involves storing a larger pool of encoded knowledge that can be dynamically grouped into adaptive local contexts, tailored to each specific query [28]. This retrieval-based strategy facilitates LTMs, which rely on subsampling, to scale more efficiently, even when the number of samples exceeds the maximum context length. Moreover, separating new knowledge from the feature space used to revisit older knowledge reduces interference between old and new subspaces, preserving valuable information that may not contribute immediately but could be useful in future distributions. To construct the most representative context at each iteration, it is essential to define a measure of coreset affinity. One approach involves designing a short-term context-wise query mechanism, where short-term memory with labels is utilized as training data to compute the loss in prediction for each coreset within the larger pool. This key-value pair-based query strategy provides a distribution matching measure for attention, which can then be used to dynamically select the top-$k$ suitable coresets based on their alignment with the current context. Alternatively, one might assume that the most relevant information for classifying a query point is contained within its local vicinity, employing k-NN to identify the nearest domain centers [142]. The number of neighbors governs the model’s expressivity and the bias-variance trade-off. Naturally, the varying potentials of coresets become particularly pronounced in practical scenarios involving imbalanced or noisy instances. To address this, a second training set can be constructed in which the sample size in each region is inversely proportional to that of the original set. This allows for the retrieval of neighboring samples from both sets for a new query [81]. When pursuing this research direction, it is essential to note that TabPFN [33] replicates each prompt across multiple permutations of feature shufflings and scalings to ensure invariance with respect to table column order. To address the resulting inefficiencies in the forward pass, TabPFN prompting fits both the test samples and their context simultaneously, thereby accelerating inference. As a result, any retrieval-based strategy built upon TabPFN must align with this batching protocol, wherein multiple test points share the same context.
5.5 Meta-representations
Rather than operating directly within the original data space, an LTM can be applied in a projected space, shifting the focus to learning how to represent relationships between features and labels in an embedding space. This approach addresses the heterogeneity within attribute and class spaces across tabular datasets, which not only complicates joint training of an LTM but also hinders its direct application during inference. To achieve this, LTMs can incorporate a meta-representation module, enabling datasets to be transformed into a uniform format of consistent dimensionality while filtering out redundant or noisy attributes. For example, TabPTM standardizes heterogeneous datasets by defining instances, regardless of their original dimensionality, through their membership to a class in terms of the Euclidean or Manhattan distance to the top-$k$ nearest class-specific prototypes in the training set [30]. This results in a meta-representation of instances as a set of $K$-dimensional vectors, one for each of the classes, which is then used to train the LTM across multiple tabular datasets to extract class-wise confidence scores. With this trained LTM, one can compute the meta-representation of any downstream dataset and perform generalization-free classification in few-shot scenarios.
5.6 Conditional computation
Rather than executing the full LTM, specific parts of the network can be activated on a per-instance basis. This strategy significantly enhances model capacity without a proportional increase in computational cost. Building on this idea, MixturePFN improves prompt effectiveness by directing new test samples to a specialized prompter context, which is fine-tuned on a cluster of training data most relevant to the query [82]. The clusters are formed by expanding and subsampling K-Means clusters, tailored to the desired context size. The effectiveness of MixturePFN prompts depends on the number of in-context prompters (ICPs), with a larger number required as the complexity and scale of data grow to capture label entropy. MixturePFN employs a tunable hyperparameter, $\gamma$, to balance efficiency and effectiveness, where the number of ICPs is given by $\lceil\gamma N_{train}/B\rceil$. Intuitively, increasing $\gamma$ enhances effectiveness but at the cost of efficiency.
5.7 Fine-tuning
While the previous strategies generally improve performance, unexpected drops may occur due to distribution shifts between the true data-generating mechanism during inference and the pretraining dataset prior [33, 122]. To address this, MixturePFN employs bootstrapping and linear adapter layers for parameter-efficient fine-tuning, minimizing the parameter count for each new expert on the assigned context, i.e., training cluster [82]. Additionally, fine-tuning has been shown to significantly improve retrieval-based training performance [83, 142]. It is important to note, however, that methods built on top of TabPFN should recognize that fine-tuning it is nontrivial, as it has been trained on a synthetic dataset prior, rather than a pretraining dataset.
Acknowledgments
This work was funded by the EU, through the Portuguese Republic’s Recovery and Resilience Plan, within the project PRODUTECH R3. It was also funded by the Portuguese Foundation for Science and Technology under project doi.org/10.54499/UIDP/00760/2020 and Ph.D. scholarship PRT/BD/154713/2023.

### v2

@@ 6. Conclusions
In the early stages of machine learning, the dominant belief was that the best way to solve a task was to meticulously design a specific dataset and tailor a model exclusively for that task. Over time, it became evident that for certain types of unstructured data, such as image recognition, a general-purpose model could be fine-tuned on specific datasets, yielding effective results. Today, we stand at a point where highly versatile models, such as general-purpose transformers, can be prompted to tackle a wide range of tasks without task-specific training. With this work, we embrace this evolving paradigm. So far, we have identified LTMs as a promising candidate for streaming solutions, warranting further research. The key challenge is not merely the size of the model or processing speed but rather the dynamic interplay between data arrival, training, recovery, and inference. As technology advances, models will naturally become smaller, faster, and more affordable. In the meantime, efforts should focus on designing algorithms that align more closely with the core principles of data stream mining: processing data in its natural order, handling labels that may be correct at the time of request but outdated or even incorrect by the time they become available, minimizing interference among concepts, and so on.
Building on this foundation, several promising research directions emerge within the main goal of context optimization in selecting data with properties similar to the desired target distribution. For instance, distribution matching does not need to be constrained to a uniform context across all query points. Instead, a more flexible approach involves storing a larger pool of encoded knowledge that can be dynamically grouped into adaptive local contexts, tailored to each specific query (bonet2024hyperfast). This retrieval-based strategy allows to go beyond the proposed temporal assumption, where recent points are likely to be of the same concept. By first applying a distribution diversification method to prioritize heterogeneity in a sample and remove redundancies, distribution matching methods can operate within a representation space that allows for similarity to be measured across data points. Thus allowing to preserve valuable information that could be useful in the future and adopt new inductive biases on how to match data distributions, such as: the smoothness assumption, where if two points reside in a high-density region are close, then so should be their corresponding outputs; the cluster assumption, where if points are in the same cluster, they are likely to be of the same concept; the manifold assumption, where high-dimensional data lies in a low-dimensional manifold. For this purpose, online synopsis techniques, such as histograms, wavelets, sketches can be used to incrementally capture the internal structure of the classes (silva2013data), and followed by an offline mining process performed on the stored micro-cluster synopsis whenever a LTM sends a request. Further visionary perspectives on these challenges and opportunities were discussed at the Streaming Continual Learning AAAI Bridge 2026 (lourencco2025bridging). The authors also plan to pursue several of these directions and welcome feedback and collaboration.
Acknowledgements.
Work funded by Portugal 2030 and the national Foundation for Science and Technology under Ph.D. scholarship PRT/BD/154713/2023, project FOXPM (COMPETE2030-FEDER-00923300 – 14731), and project doi.org/10.54499/UIDP/00760/2020.